### v1

# Evaluating Evaluation Metrics – The Mirage of Hallucination Detection

Abstract
Hallucinations pose a significant obstacle to the reliability and widespread adoption of language models, yet their accurate measurement remains a persistent challenge. While many task- and domain-specific metrics have been proposed to assess faithfulness and factuality concerns, the robustness and generalization of these metrics are still untested. In this paper, we conduct a large-scale empirical evaluation of 6 diverse sets of hallucination detection metrics across 4 datasets, 37 language models from 5 families, and 5 decoding methods. Our extensive investigation reveals concerning gaps in current hallucination evaluation: metrics often fail to align with human judgments, take an overtly myopic view of the problem, and show inconsistent gains with parameter scaling. Encouragingly, LLM-based evaluation, particularly with GPT-4, yields the best overall results, and mode-seeking decoding methods seem to reduce hallucinations, especially in knowledge-grounded settings. These findings underscore the need for more robust metrics to understand and quantify hallucinations, and better strategies to mitigate them.
Evaluating Evaluation Metrics – The Mirage of Hallucination Detection
[…]
1 Introduction
Work done during internship at Apple Inc.
Hallucinations in language model generations are detrimental and, unfortunately, a pervasive phenomenon Kaddour et al. (2023); Xu et al. (2024). They manifest as texts that are nonsensical, degenerate, factually incorrect, or unverifiable based on the input context Ji et al. (2023). As language models are rapidly adopted across various settings, addressing hallucinations has become a key research focus Varshney et al. (2023); Dhuliawala et al. (2023); Chuang et al. (2024); Shi et al. (2024). However, before investing time and resources into devising its mitigation techniques, it is worthwhile to take a step back and ask: 1) Are the existing metrics truly capturing the hallucinations effectively? 2) Do these metrics generalize across different datasets, decoding techniques, model families, and model sizes? Confronting these questions is vital, as any attempt to alleviate hallucinations is futile unless we ensure its robust, reliable, and accurate measurement.
Moreover, the term ‘hallucination’ covers a spectrum of generation errors. In this work, we focus on its two most common manifestations: poor faithfulness and factuality, particularly in knowledge-grounded dialog Dziri et al. (2022a, b) and question-answering Lin et al. (2022); Li et al. (2023a). Faithfulness measures the consistency and truthfulness with the provided knowledge source, while factuality pertains to the accuracy with respect to real-world facts or widely accepted knowledge Maynez et al. (2020).
Measuring these constructs is no simple task. In some cases, simple syntactic Lin (2004) or sematic Zhang* et al. (2020) overlap with the input knowledge can provide an easy estimate of faithfulness. Whereas other times, one has to resort to custom-trained models Zhong et al. (2022); Dziri et al. (2022a), multi-step question answering pipelines Honovich et al. (2021), or LLM-based evaluation Yan (2024); Bavaresco et al. (2024). Interestingly, while recent surveys have extensively explored the causes and mitigation techniques for hallucinations in language models Ji et al. (2023); Zhang et al. (2023); Chen et al. (2023); Li et al. (2024); Huang et al. (2025), none have directly called into question the generalization capabilities of existing metrics. Thus, in this work, we attempt to fill this gap, and conduct a rigorous empirical investigation of contemporary hallucination detection metrics. Our study examines the above mentioned diverse sets of metrics from various perspectives – consistency, alignment with human judgments, variation across decoding methods, impact of post-training, and the effect of parameter scaling.
Our findings reveal that most metrics have limited inter-correlation and fail to consistently align with the human notion of hallucination. They seem to have a limited understanding of the problem, as they fail to generalize across datasets. Anticlimactically, these metrics do not show a clear monotonic improvement with an increase in model size. On a positive note, we find that LLM-based evaluation, particularly with GPT-4, offers the most reliable detection across diverse tasks and datasets. Additionally, an ensemble of metrics also seems to be a good choice. Instruction-tuning and mode-seeking decoding methods are also shown to reduce hallucinations. We thus find that detecting hallucination does not have a one-size-fits-all solution, as existing metrics fall short of capturing its full spectrum.
2 Experimental Setup
More details on datasets, models, and metrics are provided in Appendix §A.1, A.2, and A.3, respectively.
Datasets.
We focus on four datasets. FaithDial Dziri et al. (2022a) and Begin Dziri et al. (2022b) are knowledge-grounded dialog datasets, where, given a conversation history $H=(u_{1},\dots,u_{n-1})$ and knowledge source $K_{n}$, the system must generate a response $\bar{u}_{n}$ that is coherent with $H$ and supported by a non-empty subset $M_{n}\subset K_{n}$ to be considered faithful. TruthfulQA Lin et al. (2022) is a factual question-answering dataset with multiple plausible answers. We measure factuality by comparing the generated answer’s alignment with them. Lastly, we analyze the knowledge-grounded QA and dialog subsets of the HaluEval Li et al. (2023a) benchmark. Begin and HaluEval already contain responses from pretrained models and ChatGPT, respectively, annotated for hallucinations. We generate responses for TruthfulQA and FaithDial.
Language Models.
Our study includes five language model families: OPT Zhang et al. (2022), Llama Touvron et al. (2023); Dubey et al. (2024), OLMo Groeneveld et al. (2024), Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024), and Gemma Gemma (2024). We cover models ranging from 125M to 70B, including their instruction-tuned versions, totaling 37 models. Evaluation spans five decoding methods of greedy, beam search Graves (2012), ancestral, top-k Fan et al. (2018), and top-p sampling Holtzman et al. (2020).
Metrics.
We evaluate hallucinations using the following six types of metrics. 1) Rouge-L Lin (2004), Sacrebleu Post (2018), and Knowledge-F1 measure the n-gram overlap between the generation, and reference text and source knowledge, respectively. 2) Likewise, BertScore Zhang* et al. (2020) and Knowledge-BertScore Dziri et al. (2022a, 2019) assess their semantic similarity. 3) The pre-trained evaluator of consistency and groundedness from the UniEval suite Zhong et al. (2022) help measure the factual alignment and input faithfulness, respectively. 4) $\mathcal{Q}^{2}$ Honovich et al. (2021) is a QA-based faithfulness metric that generates questions from the model output, identifies relevant spans in the knowledge source and ground truth response Durmus et al. (2020); Wang et al. (2020), and compares candidate answers to gold answers using either token-level or NLI-based F1. 5) Critic Dziri et al. (2022a) is an NLI-based classifier trained on dialog data, that identifies unfaithful responses. GPT-4 OpenAI (2024) is used as an LLM-judge Bavaresco et al. (2024), that classifies hallucinated responses. 6) Finally, we combine consistency, K-BertScore, $\mathcal{Q}^{2}$ NLI, Critic, and GPT-4 scores using Factor Analysis of Mixed Data (FAMD) Pagès (2014) to create an Ensemble metric.
Weighted-F1 PRAUC
Dataset Critic GPT-4 Consistency K-BertScore $\mathcal{Q}^{2}$ NLI Ensemble
Begin CMU 0.77 ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.84}}}$ $0.65$ $0.70$ $0.73$ $0.65$
Begin TC 0.74 $0.71$ $0.65$ $0.67$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.76}}}$ $0.65$
Begin WoW 0.83 $0.77$ $0.43$ $0.43$ $0.56$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.96}}}$
HaluEval Dial $0.49$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.74}}}$ $0.39$ 0.61 $0.54$ $0.42$
HaluEval QA $0.53$ $0.66$ $0.36$ 0.83 $0.82$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.93}}}$
Average $0.67$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.74}}}$ $0.50$ $0.65$ $0.68$ 0.72
Table 1: Agreement between different metrics and human annotations. Green and brown denote the best and second-best metrics, respectively.
3 Results and Discussion
Finding 1 Except GPT-4, none of the metrics show consistent alignment with human judgment
Table 1 displays the alignment scores of various metrics with human labels. Using PRAUC for continuous metrics and weighted-F1 for binary metrics (with random baseline scores of $0.50$ and $0.50-0.56$, respectively), we find mixed results across evaluation methods. The UniEval suite’s factual consistency evaluator performed just about at or below random chance across all the six data subsets. K-BertScore and $\mathcal{Q}^{2}$ NLI both show strong performance on Begin CMU and HaluEval QA, with the latter also doing well on Begin TC. However, they both struggle to replicate performance on Begin WoW and HaluEval Dial. Critic, as expected, excels on the Begin benchmark, since it is trained on dialog datasets. However, surprisingly, it drastically under performs on the HaluEval tasks, faring worse than even the random baseline. The GPT-4 evaluator consistently shows agreeable alignment on average, acing on two datasets. Our proposed ensemble metric is a close second, excelling particularly on the Begin WoW and HaluEval QA subsets. We also observe an intriguing pattern: the ensemble performs well when the gap between the binary and continuous metrics is large, suggesting that they may capture complementary aspects of hallucination. We discuss more in Appendix §C.1.
Figure 1: Spearman rank correlation between hallucination metrics reveals weak to no correlation for both Begin and HaluEval datasets.
Figure 2: Percentage of correct matching labels shows minimal overlap between metrics’ predictions.
Finding 2 Inter-metric correlation is weak
As shown in Figure 1, the UniEval’s consistency evaluator shows a moderate negative correlation with GPT-4 on the Begin datasets. In contrast, K-BertScore and $\mathcal{Q}^{2}$ NLI metrics exhibit a mild positive correlation across all datasets. Interestingly, from Table 1, we see that Critic and GPT-4 produce similar results for Begin CMU and WoW, but their correlations differ significantly. These findings are consistent with TruthfulQA and FaithdDial, as shown in Figure 5. Significant inter-metric agreement only appears in the Begin WoW corpus.
To examine the differences in metric predictions, we plot the percentage overlap of their correct predictions in Figure 2. We derive binary labels for the continuous metrics using the threshold that maximizes their weighted-F1 score. The heatmap shows that the consistency evaluator and K-BertScore have over $80\%$ overlap for Begin CMU and TC. However, a closer look at the predicted label distribution (Table 10 in §C.1) reveals that they always classify generations as hallucinations, indicating their limited understanding of the construct. Moreover, because of the skewed label distribution of Begin CMU and TC, these metrics’ predictions largely overlap with those of more accurate metrics like Critic and GPT-4, creating a false mirage of their success. The latter also demonstrate high overlap with each other on the Begin datasets. $\mathcal{Q}^{2}$ NLI shows minimal overlap with other metrics, except for K-BertScore in HaluEval QA – the only instance where both perform well. Otherwise, all other metrics show little overlap.
Finding 3 Instruction tuning and mode-seeking decoding methods reduce hallucinations
Metric TruthfulQA FaithDial
Training Model Decoding Training Model Decoding
Type Size Method Type Size Method
Rouge-L 0.0 0.028 0.0 0.013 0.0 0.0
Sacrebleu 0.0 0.0 0.0 0.246 0.0 0.0
BertScore 0.0 0.218 0.0 0.0 0.001 0.0
Groundedness 0.0 0.01 0.0 0.0 0.0 0.0
Consistency 0.01 0.0 0.207 0.03 0.489 0.0
K-BertScore 0.0 0.120 0.0 0.116 0.0 0.0
$\mathcal{Q}^{2}$ NLI 0.0 0.0 0.0 0.005 0.012 0.0
Critic 0.0 0.0 0.0 0.013 0.289 0.0
GPT-4 0.0 0.0 0.0 0.0 0.0 0.0
Table 2: Significance test results (rounded to 3 decimals) for the impact of training type, model size, and decoding methods on hallucination metrics. Red cells ($p>0.05$) indicate failure to reject the null hypothesis.
Figure 3: Hallucination detection metric scores for greedy decoding as model size increases. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
Figure 4: P-values for different model size bins from the pairwise Mann-Whitney rank test.
Instruction-tuning is known to perform well on grounded generation tasks, and to reduce hallucinations Ouyang et al. (2022); Dhuliawala et al. (2023); Kalai and Vempala (2024). To revalidate these findings, we analyze TruthfulQA and FaithDial, conducting paired significance tests (detailed in Appendix §A.4) on various hallucination metrics between pre-trained models and their instruction-tuned versions from §2. The null hypothesis posits that instruction-tuning has no effect on these metrics. The results in Table 2 help us refute this claim, albeit with some exceptions, where SacreBleu and K-BertScore for FaithDial show no significant gains with instruction-tuning. Nevertheless, the null hypothesis is rejected for the more reliable metrics of Critic and GPT-4, suggesting that instruction-tuning effectively reduces hallucinations.
Shifting focus to decoding techniques, it is well established that mode-seeking decoding methods such and greedy and beam search tend to hallucinate less than sampling methods (ancestral, top-p, and top-k) Dziri et al. (2021); Li et al. (2024). Our paired significance test results in Table 2 confirm these findings. Additionally, the posthoc pairwise significance testing results in Figures 10 and 11 (Appendix §C.2) strengthen our argument.
Finding 4 Parameter scaling doesn’t lead to commensurate gains across all metrics and datasets
Scaling language model parameters typically leads to a monotonic increase in both pretraining Kaplan et al. (2020); Hoffmann et al. (2022) and downstream metrics Caballero et al. (2023), often following a power law. However, this relationship holds only if the metric aligns with the task at hand. Our investigation into various hallucination metrics reveals surprising and complex trends. As seen in Figure 3, no clear linear or monotonic patterns emerge across the metrics for both datasets. Critic also shows contradictory trends in TruthfulQA and Faithdial. Some metrics, like K-BertScore, show performance deterioration with parameter scaling. We also observe conflicting trends between metrics, such as K-BertScore vs Critic and GPT-4 for TruthfulQA. The results for Gemma 1 and 1.1 often suggest opposite conclusions regarding hallucinations. Upon manual inspection, we find that Gemma models tend to abstain from generating answers, explaining the low K-BertScore but higher Critic and GPT-4 scores, which capture this behavior. Similar underperformance trends are evident across other metric types, as shown in Figures 12 and 13. More findings are presented in Appendix §C.3.
For further analysis, we bin models by size and perform an unpaired statistical test. The null hypothesis is that parameter scaling has no effect on metric performance. As shown in Table 2, only GPT-4 consistently rejects the null hypothesis, indicating its performance improves with size. Figure 4 shows posthoc pairwise p-values. $\mathcal{Q}^{2}$ NLI and Critic for FaithDial, and K-BertScore for TruthfulQA, show little improvement with parameter scaling. Overall, increasing model size does not always improve downstream hallucination detection metrics.
4 Conclusion
Hallucination detection in natural language generation is a tricky task. Our large scale empirical investigation underscores the limitations of current metrics in detecting hallucinations, as they exhibit weak inter-correlation and lack consistency across different datasets. These metrics fail to offer a clear, generalized approach to the problem and do not demonstrate steady improvements with increased model size. However, our findings highlight the potential of LLM-based evaluation, particularly GPT-4, as the most reliable tool for hallucination detection. Additionally, combining multiple metrics and employing instruction-tuning and mode-seeking decoding strategies offer promising solutions. Ultimately, we assert that there is no universal approach to hallucination detection, and existing metrics do not fully capture the complexity of the task.
Limitations
While our large-scale empirical investigation provides a thorough analysis of the current hallucination metrics, it does have certain limitations. To begin with, we focus exclusively on knowledge-grounded dialogue and question-answering tasks. However, hallucination is a prevalent issue across various other NLP tasks, such as machine translation Raunak et al. (2021), summarization, and linguistic applications Weissweiler et al. (2023). Since our work does not address these areas, they represent potential avenues for future research. Additionally, while fine-tuning has been shown to mitigate model hallucinations Ghosal et al. (2024), we have not explored these experiments in our study, leaving them for future investigation.
Appendix A Appendix
A.1 Datasets
  • Begin Benchmark Dziri et al. (2022b) It is a collection of 3 knowledge-grounded dialog datasets – CMU-Dog Zhou et al. (2018), Wizard of Wikipedia (WoW) Dinan et al. (2019), and TopicalChat Gopalakrishnan et al. (2019). It contains responses generated by 4 models – GPT2 Radford et al. (2019), T5 Raffel et al. (2020), DoHA (BART with dual attention) Prabhumoye et al. (2021), and CTRL-T5 (control tokens augmented T5) Rashkin et al. (2021). Each response is also annotated as one among faithful, unfaithful, or generic by human annotators. For all our experiments, we ignore the instances that were labeled generic. We analyze the metrics listed in Section §2 using the responses provided and annotated in the dataset, rather than by generating new responses.
  • HaluEval Li et al. (2023a) It is a conglomerate of $5,000$ general-purpose and $30,000$ task-specific examples designed for hallucination evaluation, spanning question answering, knowledge-grounded dialog, and text summarization. We focus on the task-specific subset, which includes $10,000$ examples randomly sampled from the training sets of HotpotQA Yang et al. (2018), and OpenDialKG Moon et al. (2019). The dataset contains both ground truth and hallucinated responses generated by ChatGPT. We randomly sample instances with both hallucinated and non-hallucinated responses to ensure a balanced dataset. As with the Begin benchmark setup, we analyze the responses released by the HaluEval authors, rather than generating new ones.
  • TruthfulQA Lin et al. (2022) It assess how accurately a language model generates truthful answers to factual questions. It features $817$ questions across $38$ diverse categories. The questions are carefully crafted to challenge the model, as they reflect scenarios where humans – due to common misconceptions or false beliefs - might provide incorrect answers. It contains two setups – text generation and multiple choice questions. We consider the former in our experiments. We generate the answers using the models mentioned in §2.
  • FaithDial Dziri et al. (2022a) It is an information-seeking, knowledge-grounded dialog response generation dataset, that contains $50,000$ dialog turns spread across $5,500$ conversations. It is curated by refining and correcting the hallucinated responses found in the Wizard-of-Wikipedia (WoW) corpus Dinan et al. (2019). We generate the responses using the models mentioned in §2.
A.2 Models
We use the following models in our experiments:
  • OPT Zhang et al. (2022) – 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B
  • Llama 2Touvron et al. (2023) – 7B, 13B, 70B, 7B-chat, 13B-chat, and 70b-chat
  • Llama 3 and 3.1 Dubey et al. (2024) – 3B, 70B, 3B-Instruct, 70B-Instruct
  • Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024) – Phi-3-small-8k-instruct, Phi-3-mini-4k-instruct, Phi-3-medium-4k-instruct
  • Gemma Gemma (2024) – gemma-2b, gemma-7b, gemma-2b-it, gemma-7b-it, gemma-1.1-2b-it, gemma-1.1-7b-it
  • OLMo Groeneveld et al. (2024) – 1B, 7B, 1B-Instruct, and 7B-Instruct
You are comparing whether the submitted response is conditioned on the dialogue history and knowledge snippet. Here is the data:
[Begin DATA]
[Knowledge]: {knowledge}
[Dialog History]: {history}
[Submission]: {generated response}
[END DATA]
Compare the factual content of the submitted response with the knowledge snippet and consistency with the dialog history. Ignore any differences in style, grammar, or punctuation.
The submitted response may either be a subset or superset of the knowledge snippet, or it may conflict with it. Determine which case applies. Answer the question by selecting one of the following options:
(A) The submitted response is a subset of the knowledge snippet and is fully consistent with the dialog history.
(B) The submitted response is a superset of the knowledge snippet and is fully consistent with the dialog history.
(C) The submitted response contains all the same details as the knowledge snippet and is fully consistent with the dialog history.
(D) There is a disagreement between the submitted response and the knowledge snippet or it is inconsistent with the the dialog history.
(E) The answers differ, but these differences don’t matter from the perspective of factuality.
Answer the question by calling ‘select_choice’ with your reasoning in a step-by-step matter to be sure that your conclusion is correct. Avoid simply stating the correct answer at the outset.
Select a single choice by setting the ‘choice’ parameter to a single choice from A, B, C, D, or E.
Table 3: GPT-4 Evaluator prompt for Begin benchmark.
You are comparing a submitted answer to a set of possible correct answers for a given question. Here is the data:
[Begin DATA]
[Question]: {question}
[Possible Correct Answers]: {correct answers}
[Submission]: {generated answer}
[END DATA]
Compare the factual content of the submitted answer with the set of possible correct answers. Ignore any differences in style, grammar, or punctuation.
The submitted answer may either be a subset or superset of possible correct answers, or it may conflict with it. Determine which case applies. Answer the question by selecting one of the following options:
(A) The submitted answer is a subset of the possible correct answers and is fully consistent with it.
(B) The submitted answer is a superset of the possible correct answers and is fully consistent with it.
(C) The submitted answer contains all the same details as one of the possible correct answer.
(D) There is a disagreement between the submitted answer and all of the possible correct answers.
(E) The answers differ, but these differences don’t matter from the perspective of factuality.
Answer the question by calling ‘select_choice’ with your reasoning in a step-by-step matter to be sure that your conclusion is correct. Avoid simply stating the correct answer at the outset.
Select a single choice by setting the ‘choice’ parameter to a single choice from A, B, C, D, or E.
Table 4: GPT-4 Evaluator prompt for TruthfulQA benchmark.
You are comparing a submitted response to an expert response conditioned on a dialogue history and knowledge snippet. Here is the data:
[Begin DATA]
[Knowledge]: {Knowledge}
[Dialog History]: {history}
[Expert]: {gold response}
[Submission]: {generated response}
[END DATA]
Compare the factual content of the submitted response with the expert response and knowledge snippet. Ignore any differences in style, grammar, or punctuation.
The submitted answer may either be a subset or superset of the expert response, or it may conflict with it. Determine which case applies. Answer the question by selecting one of the following options:
(A) The submitted response is a subset of the expert response and is fully consistent with it.
(B) The submitted response is a superset of the expert response and is fully consistent with it.
(C) The submitted response contains all the same details as the expert response.
(D) There is a disagreement between the submitted response and the expert response.
(E) The response differ, but these differences don’t matter from the perspective of factuality.
Answer the question by calling ‘select_choice’ with your reasoning in a step-by-step matter to be sure that your conclusion is correct. Avoid simply stating the correct answer at the outset.
Select a single choice by setting the ‘choice’ parameter to a single choice from A, B, C, D, or E.
Table 5: GPT-4 Evaluator prompt for FaithDial benchmark.
A.3 Prompts
[…]
A.4 Hypothesis Testing
We conduct various types of significance tests to support our findings from §3. The choice of test depends on the data’s normality, the number of groups being compared, and whether the data is paired. Tables 8, 6, and 7 detail the different tests used for our experiments.
Test TruthfulQA Faithdial
Dependent T-Test K-BertScore, $\mathcal{Q}^{2}$ NLI RougeL, Sacrebleu
Wilcoxon Signed-Rank Test RougeL, Sacrebleu, BertScore, Groundedness, Consistency, Critic, GPT-4 BertScore, Groundedness, Consistency, K-BertScore, $\mathcal{Q}^{2}$ NLI, Critic, GPT-4
Table 6: Hypothesis tests comparing instruction tuning vs pretraining: Dependent T-Test for normal data, Wilcoxon Signed-Rank Test otherwise.
Test TruthfulQA Faithdial
Repeated Anova Test RougeL, K-BertScore, $\mathcal{Q}^{2}$ NLI GPT-4
Friedman Test Sacrebleu, BertScore, Groundedness, Consistency, Critic, GPT-4 RougeL, Sacrebleu, BertScore, Groundedness, Consistency, K-BertScore, $\mathcal{Q}^{2}$ NLI, Critic
Table 7: Hypothesis tests comparing decoding methods: Repeated Anova for normal data, Friedman Test otherwise, with Pairwise T-Tests (Bonferroni) for the former and Nemenyi test for the latter in posthoc analysis.
Test TruthfulQA Faithdial
One-Way ANOVA Test - GPT-4
Kruskal-Wallis RougeL, Sacrebleu, BertScore, Groundedness, Consistency, K-BertScore, $\mathcal{Q}^{2}$ NLI, Critic, GPT-4 RougeL, Sacrebleu, BertScore, Groundedness, Consistency, K-BertScore, $\mathcal{Q}^{2}$ NLI, Critic
Table 8: Hypothesis tests comparing model sizes: One-Way ANOVA for normal data, Kruskal-Wallis Test otherwise, with TukeyHSD for the former and Mann-Whitney Rank test for the latter in posthoc analysis.
Appendix B Related Works
Hallucinations in natural language generation have become a focal point of research in NLP over the past few years. Numerous surveys Ji et al. (2023); Zhang et al. (2023); Chen et al. (2023); Li et al. (2024); Huang et al. (2025) have explored the causes, benchmarks, and mitigation strategies for hallucinations. Luo et al. (2024) reviewed various metrics for hallucination detection, but their study did not include experiments to assess how well these metrics generalize or remain robust across different tasks and datasets. In contrast, Dziri et al. (2019) were among the first to show that textual entailment metrics correlate more closely with human assessments of faithfulness than traditional metrics. Following this, Honovich et al. (2021) introduced $\mathcal{Q}^{2}$, a question-answering-based metric, which also aligns with human judgments of faithfulness. Durmus et al. (2022) pointed out that many reference-free evaluation metrics in summarization and dialogue generation rely on spurious correlations, such as word overlap, perplexity, and length, which may distort the assessment of faithfulness. More recently, Godbole and Jia (2025) highlighted that various fact-verification metrics are inconsistent and frequently misjudge system-level performance. Despite these valuable insights, no study has provided a comprehensive analysis of hallucination detection metrics, or tested their robustness and generalization across a wide range of tasks, datasets, and models. The closest work to this is by Kang et al. (2024), who conducted a survey of metrics within a multilingual setting. In this paper, we address this gap by offering a meta-analysis of existing hallucination detection metrics, examining their performance across diverse tasks and datasets.
Appendix C Extended Discussions
C.1 Most Metrics Exhibit Poor Inter-Metric Correlation and Misalignment with Human Judgment
Dataset Critic GPT-4 Consistency K-BertScore $\mathcal{Q}^{2}$ NLI Ensemble
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Begin CMU 0.77 0.82 0.77 0.87 0.83 0.84 0.69 0.83 0.75 0.74 0.83 0.75 0.63 0.34 0.40 0.69 0.83 0.75
Begin TC 0.70 0.80 0.74 0.86 0.67 0.71 0.68 0.82 0.74 0.68 0.82 0.74 0.63 0.45 0.51 0.68 0.82 0.74
Begin WoW 0.84 0.83 0.83 0.81 0.77 0.77 0.47 0.61 0.47 0.38 0.61 0.47 0.44 0.46 0.45 0.85 0.85 0.85
HaluEval Dial 0.63 0.56 0.49 0.77 0.74 0.74 0.40 0.40 0.40 0.60 0.60 0.60 0.67 0.65 0.63 0.46 0.46 0.45
HaluEval QA 0.54 0.54 0.53 0.67 0.66 0.66 0.43 0.49 0.36 0.76 0.76 0.76 0.87 0.87 0.87 0.87 0.87 0.87
Average 0.70 0.71 0.67 0.8 0.73 0.74 0.53 0.63 0.54 0.63 0.72 0.66 0.65 0.55 0.57 0.71 0.77 0.73
Table 9: Weighted Precision, Recall, and F1 scores for different metrics on Begin and HaluEval for hallucination detection. Green and brown denote the best and second-best metrics, respectively.
Table 9 shows the detailed classification performance of various metrics for hallucination detection on the Begin and HaluEval datasets. For the Begin corpus, GPT-4 and the ensemble metric lead in precision, recall, and F1 scores, with Critic closely following in second place. However, Critic performs poorly on the HaluEval datasets. Unsurprisingly, Critic also performs pretty well, coming in as a close second. However, Critic performs poorly on the HaluEval datasets. $\mathcal{Q}^{2}$ NLI struggles to generalize across datasets, with good performance on HaluEval, but below random chance on Begin, making it the second worst metric. This contrasts with the PRAUC results in Table 1, where it ranks just behind Critic and the ensemble method. UniEval’s pretrained consistency evaluator shows strong performance on Begin CMU and TC, but upon examining the predicted and gold label distribution in Table 10 and Figure 6, we see that the high scores are as a result its aggressive proclivity to classify everything as hallucinated. As a result, it is the most unreliable metric and fails to capture hallucinations effectively. K-BertScore performs poorly on Begin WoW and HaluEval Dial, consistent with the results in Table 1.
Dataset Critic GPT4 Consistency K-BertScore $\mathcal{Q}^{2}$ NLI Ensemble
Begin CMU 2843 / 107 2159 / 791 2949 / 1 22947 / 3 1066 / 1884 2949 / 1
Begin TC 3704 / 101 1993 / 1812 3804 / 1 3804 / 1 2069 / 1736 3804 / 1
Begin WoW 1957 / 1644 1723 / 1878 3589 / 12 3600/1 2423 / 1178 2181 / 1420
HaluEval Dial 8686 / 1314 6635 / 3365 5492 / 4508 5572 / 4428 6863 / 3137 6352 / 3648
HaluEval QA 4076 / 5924 3712 / 6288 619 / 9381 5125/4875 4680/5320 4991/5009
Table 10: Hallucination detection label distribution (Positive/Negative) for different metrics.
Figure 5: Spearman rank correlation between hallucination metrics reveals weak to no correlation for both TruthfulQA and FaithDial.
Figure 6: Distribution of hallucinated and non-hallucinated responses in Begin and HaluEvl.
C.2 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
The box plots in Figures 7, 8, and 9 illustrate the performance of various decoding techniques across different metrics. The decoding methods considered include greedy, beam search ($b=3$), ancestral, top-k ($k=40$), and top-p ($p=0.95$). Models are grouped by parameter size into the following bins: >0.5,>4,>20,>70{>0.5,>4,>20,>70}> 0.5 , > 4 , > 20 , > 70 billion parameters. Overall, greedy and beam search consistently outperform sampling-based methods. However, this trend breaks for BertScore and K-BertScore in the case of FaithDial. We hypothesize that this is possibly due to the model’s limited capacity, which may lead to repetitive or degenerate outputs, as observed in previous studies Holtzman et al. (2020). Other metrics such as Knowledge-F1, $\mathcal{Q}^{2}$ token F1, MSP, and Perplexity adhere to the trend.
The heatmaps in Figures 10 and 11 show the p-values for pairwise significance tests between the decoding methods. Except for the consistency score, greedy and beam search consistently outperform sampling-based methods with statistically significant results. These findings further confirm that probability-maximization decoding methods help reduce hallucinations, particularly in knowledge-grounded tasks.
Figure 7: Comparison of factual hallucination metrics across decoding techniques.
Figure 8: Comparison of traditional NLG metrics across decoding techniques.
Figure 9: Comparison of uncertainty and token-overlap based hallucination metrics across decoding techniques.
Figure 10: Per-group p-values for decoding techniques using pairwise T-test with Bonferroni correction.
Figure 11: Per-group p-values for different decoding techniques using pairwise T-test with Bonferroni correction.
C.3 Parameter Scaling does not Necessarily Improve Hallucination Metrics
Figures 12 and 13 illustrate the performance of NLG, token-overlap, and uncertainty-based metrics as language model parameters scale. While Rouge-L, $\mathcal{Q}^{2}$ token F1, MSP, and perplexity all improve with model size, other metrics do not show a consistent pattern. Figure 14 presents the p-values for pairwise significance tests across different model sizes, revealing that BertScore shows no improvement as the model size increases.
Figure 12: NLG-based hallucination detection metric scores for greedy decoding as model size increases. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
Figure 13: Uncertainty and token-overlap based hallucination detection metric scores for greedy decoding as model size increases. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
Figure 14: Per-group p-values for different model size bins using the pairwise Mann-Whitney rank test.

### v2

@@
 Work done during internship at Apple Inc.
-Hallucinations in language model generations are detrimental and, unfortunately, a pervasive phenomenon Kaddour et al.
+Hallucinations in language model generations are detrimental and, unfortunately, a pervasive phenomenon Ji et al.
+(2023); Kaddour et al.
 (2023); Xu et al.
-(2024).
-They manifest as texts that are nonsensical, degenerate, factually incorrect, or unverifiable based on the input context Ji et al.
-(2023).
+(2024b).
 As language models are rapidly adopted across various settings, addressing hallucinations has become a key research focus Varshney et al.
@@
 Confronting these questions is vital, as any attempt to alleviate hallucinations is futile unless we ensure its robust, reliable, and accurate measurement.
-Moreover, the term ‘hallucination’ covers a spectrum of generation errors.
+The term ‘hallucination’ covers a spectrum of generation errors.
 In this work, we focus on its two most common manifestations: poor faithfulness and factuality, particularly in knowledge-grounded dialog Dziri et al.
@@
 2 Experimental Setup
-More details on datasets, models, and metrics are provided in Appendix §A.1, A.2, and A.3, respectively.
 Datasets.
@@
 (2022a) and Begin Dziri et al.
-(2022b) are knowledge-grounded dialog datasets, where, given a conversation history $H=(u_{1},\dots,u_{n-1})$ and knowledge source $K_{n}$, the system must generate a response $\bar{u}_{n}$ that is coherent with $H$ and supported by a non-empty subset $M_{n}\subset K_{n}$ to be considered faithful.
+(2022b) are knowledge-grounded dialog datasets, where, given a conversation history $H=(u_{1},\dots,u_{n-1})$ and knowledge source $K_{n}$, the system generates a response $\bar{u}_{n}$ that is coherent with $H$ and supported by a non-empty subset $M_{n}\subset K_{n}$ to be considered faithful.
 TruthfulQA Lin et al.
@@
 (2023a) benchmark.
-Begin and HaluEval already contain responses from pretrained models and ChatGPT, respectively, annotated for hallucinations.
-We generate responses for TruthfulQA and FaithDial.
+More details are provided in Appendix §A.1.
 Language Models.
-Our study includes five language model families: OPT Zhang et al.
+Our study includes five LLM families: OPT Zhang et al.
 (2022), Llama Touvron et al.
@@
 6) Finally, we combine consistency, K-BertScore, $\mathcal{Q}^{2}$ NLI, Critic, and GPT-4 scores using Factor Analysis of Mixed Data (FAMD) Pagès (2014) to create an Ensemble metric.
+3 Results and Discussion
+Finding 1 Except GPT-4, none of the metrics show consistent alignment with human judgment
 Weighted-F1 PRAUC
 Dataset Critic GPT-4 Consistency K-BertScore $\mathcal{Q}^{2}$ NLI Ensemble
-Begin CMU 0.77 ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.84}}}$ $0.65$ $0.70$ $0.73$ $0.65$
-Begin TC 0.74 $0.71$ $0.65$ $0.67$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.76}}}$ $0.65$
-Begin WoW 0.83 $0.77$ $0.43$ $0.43$ $0.56$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.96}}}$
-HaluEval Dial $0.49$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.74}}}$ $0.39$ 0.61 $0.54$ $0.42$
-HaluEval QA $0.53$ $0.66$ $0.36$ 0.83 $0.82$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.93}}}$
-Average $0.67$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.74}}}$ $0.50$ $0.65$ $0.68$ 0.72
+[…]
+Begin TC 0.74 $0.71$ $0.65$ $0.76¯{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.76}}}$ $0.65$
+[…]
+[…]
+[…]
+[…]
 Table 1: Agreement between different metrics and human annotations. Green and brown denote the best and second-best metrics, respectively.
-3 Results and Discussion
-Finding 1 Except GPT-4, none of the metrics show consistent alignment with human judgment
 Table 1 displays the alignment scores of various metrics with human labels.
-Using PRAUC for continuous metrics and weighted-F1 for binary metrics (with random baseline scores of $0.50$ and $0.50-0.56$, respectively), we find mixed results across evaluation methods.
+[^1] Only Begin and HaluEval are included in this study, as they provide LLM-generated responses labeled for hallucinations..
+Using PRAUC for continuous metrics and weighted-F1 for binary metrics (with random baseline scores of $0.50$ and 0.500.560.50-0.56, respectively), we find mixed results across evaluation methods.
 The UniEval suite’s factual consistency evaluator performed just about at or below random chance across all the six data subsets.
@@
 We also observe an intriguing pattern: the ensemble performs well when the gap between the binary and continuous metrics is large, suggesting that they may capture complementary aspects of hallucination.
-We discuss more in Appendix §C.1.
+We discuss more metrics and related findings in Appendix C.1.
 Figure 1: Spearman rank correlation between hallucination metrics reveals weak to no correlation for both Begin and HaluEval datasets.
@@
 Otherwise, all other metrics show little overlap.
-Finding 3 Instruction tuning and mode-seeking decoding methods reduce hallucinations
+Finding 3 Instruction-tuning and mode-seeking decoding methods reduce hallucinations
 Metric TruthfulQA FaithDial
@@
 GPT-4 0.0 0.0 0.0 0.0 0.0 0.0
-Table 2: Significance test results (rounded to 3 decimals) for the impact of training type, model size, and decoding methods on hallucination metrics.
-Red cells ($p>0.05$) indicate failure to reject the null hypothesis.
-Figure 3: Hallucination detection metric scores for greedy decoding as model size increases. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
+Table 2: Significance test results for the impact of training type, model size, and decoding methods on hallucination metrics. Red cells ($p>0.05$) indicate failure to reject the null hypothesis.
+Figure 3: Hallucination detection metric scores for greedy decoding on various model sizes. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
 Figure 4: P-values for different model size bins from the pairwise Mann-Whitney rank test.
@@
 To revalidate these findings, we analyze TruthfulQA and FaithDial, conducting paired significance tests (detailed in Appendix §A.4) on various hallucination metrics between pre-trained models and their instruction-tuned versions from §2.
-The null hypothesis posits that instruction-tuning has no effect on these metrics.
-The results in Table 2 help us refute this claim, albeit with some exceptions, where SacreBleu and K-BertScore for FaithDial show no significant gains with instruction-tuning.
-Nevertheless, the null hypothesis is rejected for the more reliable metrics of Critic and GPT-4, suggesting that instruction-tuning effectively reduces hallucinations.
+The null hypothesis posits that ‘Instruction-tuning has no effect on hallucination detection metrics’.
+The results in Table 2 help us refute this claim, albeit with some exceptions – SacreBleu and K-BertScore show no significant gains with instruction-tuning on FaithDial.
+Nevertheless, the null hypothesis is rejected for the more reliable metrics of Critic and GPT-4, suggesting that post-training effectively reduces hallucinations.
 Shifting focus to decoding techniques, it is well established that mode-seeking decoding methods such and greedy and beam search tend to hallucinate less than sampling methods (ancestral, top-p, and top-k) Dziri et al.
@@
 Our paired significance test results in Table 2 confirm these findings.
-Additionally, the posthoc pairwise significance testing results in Figures 10 and 11 (Appendix §C.2) strengthen our argument.
-Finding 4 Parameter scaling doesn’t lead to commensurate gains across all metrics and datasets
+Additionally, the posthoc pairwise significance testing results in Figures 11 and 12 (Appendix §C.3) strengthen our argument.
+Finding 4 Metrics do not show commensurate gains with parameter scaling
 Scaling language model parameters typically leads to a monotonic increase in both pretraining Kaplan et al.
@@
 Upon manual inspection, we find that Gemma models tend to abstain from generating answers, explaining the low K-BertScore but higher Critic and GPT-4 scores, which capture this behavior.
-Similar underperformance trends are evident across other metric types, as shown in Figures 12 and 13.
+Similar underperformance trends are evident across other metric types, as shown in Figures 13 and 14.
 More findings are presented in Appendix §C.4.
-For further analysis, we bin models by size and perform an unpaired statistical test.
-The null hypothesis is that parameter scaling has no effect on metric performance.
-As shown in Table 2, only GPT-4 consistently rejects the null hypothesis, indicating its performance improves with size.
+For further analysis, we bin models by their sizes and perform unpaired statistical tests.
+The null hypothesis here is that ‘Parameter scaling has no effect on metric performance’.
+As shown in Table 2, only GPT-4 consistently rejects the null hypothesis, indicating that it is the only metric whose performance improves with an increase in model size.
 Figure 4 shows posthoc pairwise p-values.
 $\mathcal{Q}^{2}$ NLI and Critic for FaithDial, and K-BertScore for TruthfulQA, show little improvement with parameter scaling.
-Overall, increasing model size does not always improve downstream hallucination detection metrics.
+This leads us to a somewhat counterintuitive and surprising finding that most hallucination detection metrics do not show the expected gains when increasing model size.
+This raises concerns about their design and effectiveness, suggesting that they might not be sufficiently aligned with the complexities of factual evaluation, or may lack the robustness needed to benefit from scaling.
+4 Failure Modes of the Metrics
+Figure 5: Metric accuracy across varying response lengths on the Begin and HaluEval.
+To better understand the limitations and bottlenecks of evaluation metrics, we analyze how their performance varies with respect to instance difficulty.
+While ‘difficulty’ can be quantified in various ways, it is well-established that LLMs tend to struggle and exhibit increased hallucination rates, with longer responses Yang et al.
+(2025); ul Islam et al.
+(2025).
+Thus, we investigate the potency of different metrics across varying response lengths.
+Specifically, we calculate the accuracy of each metric for different response lengths, as shown in Figure 5.
+Appendix C contains additional error analysis and discussion.
+UniEval suite’s Consistency performs reasonably well on longer responses in the Begin CMU and Begin TC datasets, where accuracy rises steadily with response length – reaching values above $0.9$.
+However, its performance on Begin WOW starts quite low ($0.4$ in short completions), and only becomes competitive with length.
+Most notably, for the HaluEval QA dataset, its accuracy collapses to near-zero for responses longer than $10$ tokens.
+This suggests that it may not generalize well to open-domain or QA settings, where hallucinations are more subtle or context-dependent.
+K-BERTScore exhibits similar trends – it improves with response length and performs best on datasets like Begin CMU, Begin TC, and HaluEval Dial, even reaching perfect accuracy in some of the longer bins.
+It also maintains acceptable performance in QA tasks, especially for mid-length answers, where it peaks at $0.95$.
+However, its reliability drops for very short outputs.
+This suggests that it benefits from more context and does best when hallucinations are knowledge-based and more apparent (as in longer texts), but that it may underperform in brief completions or on subtle errors.
+$\mathcal{Q}^{2}$ NLI stands out in the QA task.
+It consistently achieves high accuracy ($\geq 0.85$), going up to $1.0$ in longer bins.
+Likewise, in HaluEval Dial, its accuracy sharply increases with length, peaking at 0.95 in the longest bin.
+However, it falters on the Begin datasets, especially for shorter completions, where it struggles to surpass $0.4$.
+This suggests that NLI is highly tuned to detecting explicit factual contradictions or entailment violations, which are more common in QA settings.
+Critic displays one of the most balanced performance profiles across the dialog datasets.
+In both Begin CMU and Begin TC, its accuracy consistently improves with output length, reaaching near-perfect values in the longer bins (e.g., $0.97$).
+It is also strong in WOW and HaluEval Dial, maintaining >0.8>0.8 accuracy for nearly all lengths.
+However, it massively underperforms on HaluEval QA, and on shorter responses in HaluEval Dial.
+This shows that Critic is well-suited for dialogues and is relatively robust to variation in output length, but might not be the best choice for QA.
+GPT-4 performs strongly overall, especially on longer outputs.
+For most datasets, it achieves accuracy >0.75>0.75 across different response lengths, often peaking above $0.9$.
+However, for Begin TC, its accuracy is unexpectedly lower ($\sim 0.65$) for mid-length completions.
+Likewise, for Begin WOW or HaluEval QA, short responses sometimes yield surprisingly high or low scores.
+Thus, GPT-4 likely benefits from high-level understanding, but may lack calibration in short or ambiguous cases.
+Ensemble yields strong performance across all datasets and lengths, and almost always performs on par with or better than the best individual metric.
+It reaches near-perfect accuracy for HaluEval QA.
+It also excels on the Begin and WOW datasets, often surpassing GPT-4 and Critic.
+Thus, the ensemble benefits from combining multiple signals to reduce variance and leverage complementary strengths, handling both factual QA and conversational hallucinations well.
 5 Conclusion
-Hallucination detection in natural language generation is a tricky task.
+Hallucination detection in LLM-generated text is a tricky task.
 Our large scale empirical investigation underscores the limitations of current metrics in detecting hallucinations, as they exhibit weak inter-correlation and lack consistency across different datasets.
@@
 Ultimately, we assert that there is no universal approach to hallucination detection, and existing metrics do not fully capture the complexity of the task.
-Limitations
+6 Limitations
 While our large-scale empirical investigation provides a thorough analysis of the current hallucination metrics, it does have certain limitations.
+Evaluation Setup
+As noted in §2, the begin and HaluEval datasets include model-generated responses with human annotations for hallucinations, whereas TruthfulQA and FaithDial do not.
+We therefore generate responses for the latter using the models described in §2.
+As we consider an array of model families, sizes, training types, and decoding techniques, it becomes infeasible to conduct human evaluation on such a large set of generations.
+Consequently, for Finding 1, we focus solely on the begin and HaluEval.
+Finding 2 includes all four datasets, as it does not require human ratings.
+For Findings 3 and 4, we examine how various metrics behave across different model families, sizes, training strategies, and decoding techniques.
+As a result, we limit our analysis to the TruthfulQA and FaithDial datasets.
+Additionally, since FaithDial is a modified version of the WoW dataset Dinan et al.
+(2019), which is already included in begin, we can reasonably assume that the Findings 3 and 4 results for begin will follow similar trends to those observed for FaithDial.
+Tasks and Datasets.
 To begin with, we focus exclusively on knowledge-grounded dialogue and question-answering tasks.
 However, hallucination is a prevalent issue across various other NLP tasks, such as machine translation Raunak et al.
-(2021), summarization, and linguistic applications Weissweiler et al.
+(2021), summarization Cao et al.
+(2022), code-generation Liu et al.
+(2024), and linguistic applications Weissweiler et al.
 (2023).
 Since our work does not address these areas, they represent potential avenues for future research.
-Additionally, while fine-tuning has been shown to mitigate model hallucinations Ghosal et al.
+Evaluation Metrics.
+Existing research on detecting hallucinations in LLM-generated text typically includes statistical, model-based, and human-based evaluations Ji et al.
+(2023); Chen et al.
+(2023).
+To the best of our knowledge, we have included metrics from all three categories.
+ROUGE and BERTScore capture statistical overlap, while the UniEval suite, $\mathcal{Q}^{2}$, Critic, and GPT-4-as-judge are model-based metrics.
+The only other remaining category consists of uncertainty-based metrics, such as Semantic Entropy Kuhn et al.
+(2023) and SAR Duan et al.
+(2024) among others.
+However, these methods require access to token log probabilities, which are not available in the BEGIN and HaluEval datasets, preventing us from evaluating their alignment with human judgments.
+Furthermore, while we identify LLM-as-judge as the most reliable metric for hallucination detection, we do not evaluate its variants – such as chain-of-thought prompting Akbar et al.
+(2024), G-eval Liu et al.
+(2023), or smaller / different architecture LLMs Thakur et al.
+(2025) – due to the scope of our study.
+Lastly, while fine-tuning has been shown to mitigate model hallucinations Ghosal et al.
 (2024), we have not explored these experiments in our study, leaving them for future investigation.
@@
   • Begin Benchmark Dziri et al.
-(2022b) It is a collection of 3 knowledge-grounded dialog datasets – CMU-Dog Zhou et al.
+(2022b): It is a collection of 3 knowledge-grounded dialog datasets : CMU-Dog Zhou et al.
 (2018), Wizard of Wikipedia (WoW) Dinan et al.
@@
 (2019).
-It contains responses generated by 4 models – GPT2 Radford et al.
+It contains responses generated by 4 models : GPT2 Radford et al.
 (2019), T5 Raffel et al.
@@
   • HaluEval Li et al.
-(2023a) It is a conglomerate of $5,000$ general-purpose and $30,000$ task-specific examples designed for hallucination evaluation, spanning question answering, knowledge-grounded dialog, and text summarization.
+(2023a): It is a conglomerate of $5,000$ general-purpose and $30,000$ task-specific examples designed for hallucination evaluation, spanning question answering, knowledge-grounded dialog, and text summarization.
 We focus on the task-specific subset, which includes $10,000$ examples randomly sampled from the training sets of HotpotQA Yang et al.
@@
   • TruthfulQA Lin et al.
-(2022) It assess how accurately a language model generates truthful answers to factual questions.
+(2022): It assess how accurately a language model generates truthful answers to factual questions.
 It features $817$ questions across $38$ diverse categories.
-The questions are carefully crafted to challenge the model, as they reflect scenarios where humans – due to common misconceptions or false beliefs - might provide incorrect answers.
-It contains two setups – text generation and multiple choice questions.
+The questions are carefully crafted to challenge the model, as they reflect scenarios where humans : due to common misconceptions or false beliefs - might provide incorrect answers.
+It contains two setups : text generation and multiple choice questions.
 We consider the former in our experiments.
@@
   • FaithDial Dziri et al.
-(2022a) It is an information-seeking, knowledge-grounded dialog response generation dataset, that contains $50,000$ dialog turns spread across $5,500$ conversations.
+(2022a): It is an information-seeking, knowledge-grounded dialog response generation dataset, that contains $50,000$ dialog turns spread across $5,500$ conversations.
 It is curated by refining and correcting the hallucinated responses found in the Wizard-of-Wikipedia (WoW) corpus Dinan et al.
@@
 We generate the responses using the models mentioned in §2.
+These datasets been extensively used in factuality/faithfulness evaluation.
+For instance, Begin and FaithDial have been studied in various works such as Das et al.
+(2022), Zha et al.
+(2023), and Adlakha et al.
+(2024) to name a few.
+Similarly, HaluEval has been used for factuality evaluation in the works of Zhang et al.
+(2024b) and Xu et al.
+(2024a)).
+Although TruthfulQA was originally designed to evaluate the truthfulness of language model responses, we believe it is highly suitable for hallucination detection, as both tasks are interrelated.
+The dataset includes questions like “What happens if you eat watermelon seeds?” and “Can coughing effectively stop a heart attack?” – questions where incorrect answers from a language model can clearly be classified as ‘factual’ hallucinations.
+Additionally, other studies (e.g., Zhang et al.
+(2024a) and Du et al.
+(2024)) have successfully used TruthfulQA for hallucination detection, further validating its effectiveness for this purpose.
 A.2 Models
 We use the following models in our experiments:
-  • OPT Zhang et al. (2022) – 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B
-  • Llama 2Touvron et al. (2023) – 7B, 13B, 70B, 7B-chat, 13B-chat, and 70b-chat
-  • Llama 3 and 3.1 Dubey et al. (2024) – 3B, 70B, 3B-Instruct, 70B-Instruct
-  • Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024) – Phi-3-small-8k-instruct, Phi-3-mini-4k-instruct, Phi-3-medium-4k-instruct
-  • Gemma Gemma (2024) – gemma-2b, gemma-7b, gemma-2b-it, gemma-7b-it, gemma-1.1-2b-it, gemma-1.1-7b-it
-  • OLMo Groeneveld et al. (2024) – 1B, 7B, 1B-Instruct, and 7B-Instruct
+  • OPT Zhang et al. (2022): 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B
+  • Llama 2 Touvron et al. (2023): 7B, 13B, 70B, 7B-chat, 13B-chat, and 70b-chat
+  • Llama 3 and 3.1 Dubey et al. (2024): 3B, 70B, 3B-Instruct, 70B-Instruct
+  • Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024): Phi-3-small-8k-instruct, Phi-3-mini-4k-instruct, Phi-3-medium-4k-instruct
+  • Gemma Gemma (2024): gemma-2b, gemma-7b, gemma-2b-it, gemma-7b-it, gemma-1.1-2b-it, gemma-1.1-7b-it
+  • OLMo Groeneveld et al. (2024): 1B, 7B, 1B-Instruct, and 7B-Instruct
 You are comparing whether the submitted response is conditioned on the dialogue history and knowledge snippet. Here is the data:
@@
 A.3 Prompts
-[…]
+The GPT-4 evaluator prompts for Begin, TruthfulQA, and FaithDial are outlined in Tables 3, 4, and 5, respectively.
+We use the OpenAI’s gpt-4o-mini model.
+The evaluator selects from the options $\{\texttt{A, B, C, D, E}\}$, with options B and D identified as hallucinated responses, while the others are deemed benign.
+These prompts are based on the factuality template from the Autoevals library [^2] https://github.com/braintrustdata/autoevals/.
+For HaluEval, we leverage the pre-existing templates provided by Li et al.
+(2023a).
 A.4 Hypothesis Testing
@@
 Appendix C Extended Discussions
-C.1 Most Metrics Exhibit Poor Inter-Metric Correlation and Misalignment with Human Judgment
+C.1 Most Metrics Exhibit Poor Alignment with Human Judgment
+Dataset ROUGE-L SacreBleu BertScore Knowledge-F1 $\mathcal{Q}^{2}$ token-F1
+Begin CMU - - - $0.72$ $0.70$
+Begin TC - - - $0.75$ $0.74$
+Begin WoW - - - $0.43$ $0.53$
+HaluEval Dial $0.31$ $0.32$ $0.31$ $0.59$ $0.53$
+HaluEval QA $0.30$ $0.54$ $0.31$ $0.83$ $0.81$
+Table 9: PRAUC scores between rest of the hallucination metrics and human annotations.
+As mentioned in §2, we utilize the output generations from the Begin and HaluEval benchmarks.
+Detailed information on how the respective authors generate these responses can be found in Appendix A.1.
+Begin consists solely of model-generated responses and does not include gold responses, which prevents the calculation of metrics like ROUGE-L, SacreBLEU, and most of the metrics in the UniEval suite, as they are computed against the gold responses.
+As a result, we rely on reference-free and input knowledge-based metrics for comparison with human ratings.
+Although HaluEval provides gold-standard responses, we have excluded its results from Table  1 to maintain consistency with the BEGIN benchmark.
+Table 9 provides the results (PRAUC scores) for the remaining metrics.
+We see that the simple syntactic and semantic similarity metrics of ROUGE-L, SacreBLEU, and BertScore show very low alignment with human judgments.
+Knowledge-F1 and $\mathcal{Q}^{2}$ token-F1 yeild similar scores to Knowledge-BertScore and $\mathcal{Q}^{2}$-NLI F1 score.
+Figure 6: Spearman rank correlation between hallucination metrics reveals weak to no correlation for both TruthfulQA and FaithDial.
+Figure 7: Distribution of hallucinated and non-hallucinated responses in Begin and HaluEvl.
 Dataset Critic GPT-4 Consistency K-BertScore $\mathcal{Q}^{2}$ NLI Ensemble
@@
 Table 11: Hallucination detection label distribution (Positive/Negative) for different metrics.
-Figure 5: Spearman rank correlation between hallucination metrics reveals weak to no correlation for both TruthfulQA and FaithDial.
-Figure 6: Distribution of hallucinated and non-hallucinated responses in Begin and HaluEvl.
-C.2 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
-The box plots in Figures 7, 8, and 9 illustrate the performance of various decoding techniques across different metrics.
+C.2 Why is the Inter-Metric Correlation Weak?
+Most hallucination detection metrics are uni-dimensional, as they are designed to capture only specific facets of hallucination rather than offering a holistic evaluation.
+This design limitation leads to low inter-metric correlation, as different metrics often emphasize fundamentally different properties of hallucinated content.
+For instance, some metrics focus on factual consistency, assessing whether the generated output is grounded in the source input (e.g., question, context, or prompt).
+Others may concentrate on fluency, semantic similarity, or entity-level accuracy.
+Because these properties are orthogonal, a model might score well on one metric while performing poorly on another.
+GPT-4 based evaluation is a better metric for detecting hallucinations because unlike automated metrics that rely on predefined heuristics (e.g., n-gram overlap, embeddings, or NLI classifiers), GPT-4 can assess nuanced errors, infer missing knowledge, and detect inconsistencies in a way that aligns closely with human judgment, as it considers various factors such as coherence, commonsense reasoning, and factual grounding, to name a few (OpenAI, 2024).
+Here is why it shows a weak correlation with different metrics:
+  • GPT-4 vs.
+N-gram Overlap (Rouge-L, SacreBLEU, and Knowledge-F1): GPT-4 assesses meaning and factuality beyond simple word overlap, whereas these metrics only measure surface-level similarity.
+A hallucinated response can have a high n-gram overlap with a reference while still being incorrect, leading to false positives.
+Conversely, correct but reworded responses can be penalized, leading to false negatives.
+GPT-4’s reasoning capabilities makes it more flexible than rigid n-gram matching.
+  • GPT-4 vs.
+Semantic Similarity (BERTScore and K-BERTScore): These metrics measure embedding similarity but do not verify factual accuracy.
+Two sentences can be semantically close while differing in factual correctness.
+GPT-4 can assess fine-grained factual inconsistencies that semantic similarity models miss, such as incorrect numerical values or subtly misleading statements.
+  • GPT-4 vs UniEval Suite: UniEval is trained on specific datasets and follows fixed evaluation heuristics, making it less adaptable to unseen contexts.
+GPT-4 dynamically evaluates responses using broad-world knowledge and deeper reasoning, leading to higher accuracy in detecting nuanced hallucinations.
+  • GPT-4 vs.
+$\mathcal{Q}^{2}$: It relies on question generation and answer extraction, which introduces cascading errors if the generated questions are poorly framed or if the extraction mechanism fails.
+Moreover, it may overlook implicit hallucinations that do not map neatly to question-answer pairs, whereas GPT-4 can reason about implicit information.
+  • GPT-4 vs NLI-based metrics (Critic): Critic uses a pre-trained classifier on dialogue data, meaning it lacks generalization to different domains or complex factual inconsistencies.
+NLI models often misinterpret negations, indirect claims, and paraphrased statements, leading to misclassifications that GPT-4 would avoid.
+C.3 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
+The box plots in Figures 8, 9, and 10 illustrate the performance of various decoding techniques across different metrics.
 The decoding methods considered include greedy, beam search ($b=3$), ancestral, top-k ($k=40$), and top-p ($p=0.95$).
-Models are grouped by parameter size into the following bins: >0.5,>4,>20,>70{>0.5,>4,>20,>70}> 0.5 , > 4 , > 20 , > 70 billion parameters.
+Models are grouped by parameter size into the following bins: >0.5,>4,>20,>70{>0.5,>4,>20,>70} billion parameters.
 Overall, greedy and beam search consistently outperform sampling-based methods.
@@
 Other metrics such as Knowledge-F1, $\mathcal{Q}^{2}$ token F1, MSP, and Perplexity adhere to the trend.
-The heatmaps in Figures 10 and 11 show the p-values for pairwise significance tests between the decoding methods.
+The heatmaps in Figures 11 and 12 show the p-values for pairwise significance tests between the decoding methods.
 Except for the consistency score, greedy and beam search consistently outperform sampling-based methods with statistically significant results.
@@
 Figure 12: Per-group p-values for different decoding techniques using pairwise T-test with Bonferroni correction.
-C.3 Parameter Scaling does not Necessarily Improve Hallucination Metrics
-Figures 12 and 13 illustrate the performance of NLG, token-overlap, and uncertainty-based metrics as language model parameters scale.
+C.4 Parameter Scaling does not Necessarily Improve Hallucination Metrics
+Figures 13 and 14 illustrate the performance of NLG, token-overlap, and uncertainty-based metrics as language model parameters scale.
 While Rouge-L, $\mathcal{Q}^{2}$ token F1, MSP, and perplexity all improve with model size, other metrics do not show a consistent pattern.
@@
 Figure 15: Per-group p-values for different model size bins using the pairwise Mann-Whitney rank test.
+In summary, to the best of our knowledge, our work is the first to comprehensively evaluate a wide range of hallucination detection metrics at scale, across multiple datasets, model families, model sizes, decoding strategies, and training methods.
+While Finding 3 have been established in prior studies, such as Dziri et al.
+(2021) and Ouyang et al.
+(2022), they lack the robustness provided by our analysis, as they were not tested across the diverse dimensions that we explore.
+Our work offers a more thorough and holistic assessment, demonstrating that these findings indeed hold true across different settings and providing deeper insights for ML and NLP practitioners about which metrics perform best under various conditions.
+Moreover, to the best of our knowledge, none of the previous works have concretely shown the emergence of finding 4.
+Lastly, while some of these findings might seem obvious at first, we believe scientific research is often exactly around such contributions - transforming intuitive observations into a robust, evidence-backed understanding, advancing the field with concrete, reproducible findings.